A. Giving a test to a group of students, the grades and gender are summarized below

A B C Total
Male 18 3 2 23
Female 5 10 13 28
Total 23 13 15 51


Let pi represent the percentage of all male students who would receive a grade of B on this test. Use a 90% confidence interval to estimate ππ to three decimal places.
Enter your answer as a tri-linear inequality using decimals (not percents).___<pi<___

B. Out of 400 people sampled, 104 had kids. Based on this, construct a 95% confidence interval for the true population proportion of people with kids.
Give your answers as decimals, to three places__<__
Answer #1

Ans:

A)

sample proportion=3/23=0.1304

90% confidence interval for p

=0.1304+/-1.645*sqrt(0.1304*(1-0.1304)/23)

=0.1304+/-0.1155

=(0.015, 0.246)

0.015<p<0.246

B)sample proportion=104/400=0.260

95% confidence interval for p

=0.260+/-1.96*sqrt(0.260*(1-0.260)/400)

=0.260+/-0.043

=(0.217, 0.303)

0.217<p<0.303